# SALT: A Flexible Semi-Automatic Labeling Tool for General LiDAR Point Clouds with Cross-Scene Adaptability and 4D Consistency

### v1

@@ 4.3.1 Better Views Yield Closer Textures
4.4 Tool Discussion
Our success in zero-shot segmentation has laid a solid foundation for constructing annotation tools. We compared three representative open-source LiDAR annotation tools in Tab. 6. Our method outperforms existing tools in terms of segmentation performance, tracking quality, visualization, and operation. More importantly, the operation of presegmentation does not require manual interaction, and it can generates the reference for the entire sequence. In addition, the presegmentation methods of other approaches make it difficult to achieve zero-shot segmentation across different types of LiDAR.
@@ 5 Conclusion
Future improvements can focus on two key aspects. One is refining our data alignment framework by replacing SAM2 with a novel VFM that offers superior tracking performance. Another is addressing the sensitivity of the greedy algorithm to initial values in data alignment. We anticipate that open-sourcing of SALT will scale up LiDAR data accessibility, driving exponential growth in LiDAR datasets and laying a foundation for future LiDAR foundation models.
[…]
Acknowledgments This work is supported by the National Key R&D Program of China (Grant 2020YFC2007500), and the Science and Technology Commission of Shanghai Municipality (Grant 20DZ2220400). Yanbo Wang and Yongtao Chen contributed equally to this work. Jingchuan Wang (jchwang@sjtu.edu.cn) and Weidong Chen (wdchen@sjtu.edu.cn) are the corresponding authors.
[…]
Data Availability Our code is available in https://github.com/Cavendish518/SALT. Datasets are available as follows: S.MID [65] in https://semanticindustry.top/, SemanticKITTI [4] in https://semantic-kitti.org/, SemanticKITTI-16 [76] in https://yanx27.github.io/RobustLidarSeg/, and nuScenes [7] in https://www.nuscenes.org/.
Appendix A User Manual of SALT
The success of LabelMe [57] has made us aware of the significant contribution that a user-friendly annotation tool can make to the prosperity of the community. We have integrated a fully automatic annotation module into LABELER [4], as shown in Fig. 8, Fig. 9 and Fig. 10. An illustration video can be found in the project video.
@@ Appendix A User Manual of SALT
A.1 Automatic Segmentation of Whole Sequence
After loading the raw point cloud sequence data, the user only needs to click the “SALT” button once to obtain the presegmentation results for the entire sequence as shown in Fig. 8. Once the progress bar, which represents the zero-shot segmentation algorithm described in the main text, is complete, the presegmentation results are automatically saved for subsequent semantic and instance labeling. The presegmentation results are also automatically displayed in the user interface with different colors.
@@ A.1 Automatic Segmentation of Whole Sequence
A.2 Manual Assignment and Refinement for Semantic Annotation
The user is free to define as many semantic classes appear in the sequence. Users can assign custom semantic labels to the presegmentation results based on their needs as shown in Fig. 9. By simply clicking on a predefined color button representing a specific semantic category and then selecting a point cloud with a particular ID, all points with that ID will be assigned to the chosen label and updated to the corresponding color. This operation is as intuitive and effortless as a coloring game. Please note that the colors used to display the presegmentation results are designed to avoid conflict with user-defined semantic label colors. If users are not satisfied with the pre-annotated results, they can modify them using the polygon tool. Inherited from LABELER, SALT supports the option to hide other classes, making manual annotation adjustments more convenient.
@@ A.2 Manual Assignment and Refinement for Semantic Annotation
A.3 Auto Ordering and Manual Refinement for Instance Annotation
Once users are satisfied with the semantic labeling results, they can simply click the “Auto Instance” button to automatically assign instance IDs to all semantic categories as shown in Fig. 10. Users can then further refine the results by splitting or merging instance IDs within each category.
Appendix B Library of SALT
B.1 Projection for Pseudo-Image
LiDAR Point clouds are inherently sparse, and directly projecting the original point cloud into an pseudo-image typically results in a lack of continuity. A straightforward approach is to voxelize the point cloud and fill the voxels with point data for projection. Assume that each voxel is filled with $m$ points (a parameter related to the pseudo-camera’s intrinsic parameters). In general, $m>10^{6}$, and the time complexity for rendering an image in this manner is $O(mv)$, where $v$ represents the number of voxels. Considering the characteristics of LiDAR data, we proposes an accelerated image rendering technique. For each voxel, only the 8 corner points are considered. These corner points are projected onto the image plane, and we construct a convex hull. The pixels within the convex hull are then rendered. The time complexity of this method is $O(8v)$, significantly reducing rendering time.
B.2 Unprojection for Presegmentation Results
Point cloud growth is employed to reconstruct the results after segmented and tracked by SAM2, while 3D NMS serves as the foundation for the 4D NMS introduced in the main text. The corresponding algorithm is shown in Algorithm. 2.
@@ B.2 Unprojection for Presegmentation Results
1:Input: $L_{object}$, Mask
2:Output: $Y_{object}^{3dnms}$
3:Step 1: Unprojection based on the given mask.
4:Partial_segmented_pointcloud $\leftarrow$ Unproject ($L_{object}$, Mask)
5:Partial_segmented_voxels $\leftarrow$ Mapping (Partial_segmented_pointcloud)
6:Step 2: Perform region growth to get segmented voxel clusters.
[…]
8:Step 3: Process each partial segmented clusters individually.
9:for each Cluster in Partial_segmented_clusters do
10:     Step 3.1: Refine cluster’s labels using the reduce bleeding operation.
11:     Refined_voxels $\leftarrow$ Nerf_bleeding (Cluster)
12:     Step 3.2: Perform 3D NMS based on bounding boxes.
13:     Refined_voxels $\leftarrow$ NMS3d (Refined_voxels)
14:     Step 3.3: Expand labels to unlabeled voxels, prioritizing the most frequent label.
15:     Refined_voxels $\leftarrow$ Label_growth (Refined_voxels)
16:end for
17:Step 4: Convert the refined voxel representation into final labels.
18:$Y_{object}^{3dnms}$leftarrow Voxel_to_label (Refined_voxels)
B.3 Parallelizing SAM2 Inference for Efficiency
Given that SAM2 exhibits a relatively low frame rate, we propose a solution by leveraging multi-process parallelism to accelerate its inference. We perform asynchronous inference for each prompt, while also maximizing the usage of the GPU memory to ensure that computational units are fully utilized. This simple approach significantly reduces the processing time.

### v2

@@ IV-C1 Better Views Yield Closer Textures
IV-D Tool Discussion
Our success in zero-shot segmentation has laid a solid foundation for constructing annotation tools. We compared three representative open-source LiDAR annotation tools in Tab. VI. Our method outperforms existing tools in terms of segmentation performance, tracking quality, visualization, and operation. More importantly, the operation of presegmentation does not require manual interaction, and it can generates the reference for the entire sequence. In addition, the presegmentation methods of other approaches make it difficult to achieve zero-shot segmentation across different types of LiDAR.
@@ V Conclusion
Future improvements can focus on two key aspects. One is refining our data alignment framework by replacing SAM2 with a novel VFM that offers superior tracking performance. Another is addressing the sensitivity of the greedy algorithm to initial values in data alignment. We anticipate that open-sourcing of SALT will scale up LiDAR data accessibility, driving exponential growth in LiDAR datasets and laying a foundation for future LiDAR foundation models. By enabling more accurate and scalable LiDAR perception, these improvements will empower robots to perceive and interact with the physical world in a more robust and intelligent manner.
Appendix A: User Manual of SALT
The success of annotation tools [88] for 2D image has made us aware of the significant contribution that a user-friendly annotation tool can make to the prosperity of the community. We have integrated a fully automatic annotation module into LABELER [15], as shown in Fig. 8, Fig. 9 and Fig. 10. An illustration video can be found in the project video.
@@
Automatic Segmentation of Whole Sequence. After loading the raw point cloud sequence data, the user only needs to click the “SALT” button once to obtain the presegmentation results for the entire sequence as shown in Fig. 8. Once the progress bar, which represents the zero-shot segmentation algorithm described in the main text, is complete, the presegmentation results are automatically saved for subsequent semantic and instance labeling. The presegmentation results are also automatically displayed in the user interface with different colors.
@@
Manual Assignment and Refinement for Semantic Annotation. The user is free to define as many semantic classes appear in the sequence. Users can assign custom semantic labels to the presegmentation results based on their needs as shown in Fig. 9. By simply clicking on a predefined color button representing a specific semantic category and then selecting a point cloud with a particular ID, all points with that ID will be assigned to the chosen label and updated to the corresponding color. This operation is as intuitive and effortless as a coloring game. Please note that the colors used to display the presegmentation results are designed to avoid conflict with user-defined semantic label colors. If users are not satisfied with the pre-annotated results, they can modify them using the polygon tool. Inherited from LABELER, SALT supports the option to hide other classes, making manual annotation adjustments more convenient.
@@
Auto Ordering and Manual Refinement for Instance Annotation. Once users are satisfied with the semantic labeling results, they can simply click the “Auto Instance” button to automatically assign instance IDs to all semantic categories as shown in Fig. 10. Users can then further refine the results by splitting or merging instance IDs within each category.
Appendix B: Library of SALT
Projection for Pseudo-Image. LiDAR Point clouds are inherently sparse, and directly projecting the original point cloud into an pseudo-image typically results in a lack of continuity. A straightforward approach is to voxelize the point cloud and fill the voxels with point data for projection. Assume that each voxel is filled with $m$ points (a parameter related to the pseudo-camera’s intrinsic parameters). In general, $m>10^{6}$, and the time complexity for rendering an image in this manner is $O(mv)$, where $v$ represents the number of voxels. Considering the characteristics of LiDAR data, we proposes an accelerated image rendering technique. For each voxel, only the 8 corner points are considered. These corner points are projected onto the image plane, and we construct a convex hull. The pixels within the convex hull are then rendered. The time complexity of this method is $O(8v)$, significantly reducing rendering time.
Unprojection for Presegmentation Results. Point cloud growth is employed to reconstruct the results after segmented and tracked by SAM2, while 3D NMS serves as the foundation for the 4D NMS introduced in the main text. The corresponding algorithm is shown in Algorithm. 2.
@@
  Input: $L_{object}$, Mask
  Output: $Y_{object}^{3dnms}$
  Step 1: Unprojection based on the given mask.
  Partial_segmented_pointcloud $\leftarrow$ Unproject($L_{object}$, Mask)
  Partial_segmented_voxels $\leftarrow$ Mapping(Partial_segmented _pointcloud)
  Step 2: Perform region growth to get segmented voxel clusters.
  Partial_segmented_clusters $\leftarrow$ Region_growth(Partial _segmented_voxels)
  Step 3: Process each partial segmented cluster individually.
  for each Cluster in Partial_segmented_clusters do
     Step 3.1: Refine cluster’s labels using the reduce bleeding operation.
     Refined_voxels $\leftarrow$ Nerf_bleeding(Cluster)
     Step 3.2: Perform 3D NMS based on bounding boxes.
     Refined_voxels $\leftarrow$ NMS3d (Refined_voxels)
     Step 3.3: Expand labels to unlabeled voxels, prioritizing the most frequent label.
     Refined_voxels $\leftarrow$ Label_growth (Refined_voxels)
  end for
  Step 4: Convert the refined voxel representation into final labels.
  $Y_{object}^{3dnms}$leftarrow Voxel_to_label (Refined_voxels)
Parallelizing SAM2 Inference for Efficiency. Given that SAM2 exhibits a relatively low frame rate, we propose a solution by leveraging multi-process parallelism to accelerate its inference. We perform asynchronous inference for each prompt, while also maximizing the usage of the GPU memory to ensure that computational units are fully utilized. This simple approach significantly reduces the processing time.